OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - 2y = 4 4x + 5y = -18 Answers: 4x + 5(-4x - 2) = -18 4x + 5(4x - 2) = -18 4(-4x - 2) + 5y = -18 4(4x - 2) + 5y = -18 (my answer is B)

OpenStudy (anonymous):

Looks sound

OpenStudy (anonymous):

solve for y in the first equation -2y = 4-8x y = 4x-2 substitute into 2nd equation 4x + 5(4x-2) = -18 answer is B your second answer